FEL Upgrade Project Weekly Brief –June 17-21, 2002




The last major procurement for the IR Upgrade, the large vacuum chambers for the recirculation arcs went out to bid this week. The measurements on the test stand for the last of the quadrupole magnets were completed this week. Thus all the quads required for both the IR and UV upgrades are now available for installation.


The Upgrade Project monthly report for May 2002 was completed this week and distributed to the DOE, ONR and AFRL contract monitors. Both the IR and UV Upgrade projects are on track; however, because JLab is still due $1.5M of AF funding and $1.2M of ONR/JTO funding for FY02, there are budget and cash flow problems building as we approach the end of the fiscal year. We have asked for the assistance of several folks in our supply chain to help solve these funding release problems.

On Tuesday, June 18th, G. Neil attended a meeting of the various groups that are participating in

in laser lethality studies funded by the JTO. The meeting was held at the Navy’s Dahlgren Labs.

Our plan for offering the FEL User facility as a nationwide resource for continuing JTO (and other DOD funded) lethality studies is gaining support. We plan on continuing our partnership with NRL and MIT Lincoln Labs with these studies.

On Tuesday, June 18th, F. Dylla and G. Williams attended a meeting at UVA to continue our discussions on use of the FEL for materials science and biology/biomedical studies and planning for a collaboration in the use of advanced light sources.

On Thursday, June 20th, we had visits from Major Steve Redifer, who will become the Marine Corp. representative on directed energy development. Major Redifer is currently finishing graduate studies at the NPS with our colleague, Prof. Bill Colson. Also on Thursday we had a meeting with Alan Todd and Tony Favale from AES, Inc. to discuss our progress on the CRADA

for the development of the 100 mA injector.

On Friday, F.Dylla and G.Neil will be attending the next meeting of the JTO FEL "Technical Working Group" chaired by Michael Dietchman from ONR which is being hosted by the Univ. of MD (c/o Prof. Pat O’Shea). The working group is continuing an activity begun at the May meeting to revise and prioritize the technology development plan for FELs for defense applications.

WBS 4 (Injector):

The HEPA filter system for the W&M implanter is running and additional CO2 cleaning will be used on the next test tube. The ball was polished to completion, needs cleaning. Detailing of the BPM/shielded bellows is being done.

Gun HVPS - The Tank for the Gun's HVPS is being tested for leaks today. The test is progressing very well and the tank is expected to be delivered this coming Monday. New resistive material is on order and expected next week for the next shot at producing the conditioning resistor.

WBS 6 (RF):

Quarter Klystrons - The klystrons and their solenoids for the Quarter were installed on their roll-around carts this week. The plumbing and wiring will be finished as time permits.

Quarter HVPS - The PLC's for cavity 3 were installed this week. The preliminary testing of the new PLC and code has started. The PLC's and wiring for cavity 4 will be completed as time permits.

Zone 4 - The test computer and program were reactivated to finish testing the klystron and circulator for position 4-8. The second FEL cryomodule (returning from SL21in CEBAF) is expected to be installed on July 9th with a cooldown on July 24th.

WBS 8 (Instrumentation):

Progress has been made on the drive laser pulse control system. The design has progressed to where the ConOptics Model 305 (Synchronous Countdown Chassis which we have used for micropulse control) eliminated by placing the core logic board on to a VME carrier for better integration into the control system. The 6U VME board is being designed to support & power the board. The motivation is to package the entire system in a single crate and eliminate the many cables and interfaces to the various laser control chassis. Design work also continues on a board that monitors and interlocks the external RF to the drive laser. The external oscillator for the mode locker will also be mounted to a VME board.

The BLM rear cards were sent out last week and are due to ship back to us on Monday. All parts are (including front panels) are on order). The fabrication of the Beam Loss Monitor assemblies has been completed. Many thanks to our new student Nate Johnson! Electrical testing and Q.A. of the assemblies has begun. Dark current and gain measurements of the PMT's will begin next week.

A meeting was held to discuss the requirements for the 100 amp sextapole/octapole power supplies. The issues were whether we could take credit for damping of the ripple voltage since they are solid core magnets. The plan is to move forward with a low cost power supply and purchase an external current transductor. The piece price of the power supplies is ~2500 + 1000 for the transductor vs ~$15k for a 10PPM supply. If the resulting magnetic field ripple is too high (as measured on test stand) the supplies will be fitted with the current transductor and current loop regulator with sufficient gain at frequencies 60 Hz etc. The controls for this approach will rely on same devices as for the super trim supplies. This is the SNS/BNL fiber isolated controller (EPICS software is already complete!).

The front panel pcb schematic for the OMS chassis (8 channel stepper motor driver) has been completed and submitted to EECAD. The fast raster control board design is complete and submitted to EECAD, this is used to create the recirculating dump raster pattern. Power cables for the first 32 channels of trim magnets for the Injector (MFF0F01 thru MQX1F03) have been pulled into place. The rework of the injector cables continues.

Efforts continue to clean up the W&M ion implantation system. With the help of the students the RG58 cables have been replaced with .141 fully shielded super flex. Also a remote control box is being built to replace the DG535 timing generator and to allow for the rack door to be closed (shielding the pulse forming network) while bringing up to high voltage pulser.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• Magnet Enterprises International (MEI) in Oakland CA. baked out Coil #12 at 270°F (rather than 130) for 16 hours (rather than 4) and degassed for 4 to 6 hours (rather than 1.5) This did not fix the welt problem. The epoxy formulator is sending MEI s similar epoxy with out flexiblizer. We will pot # 13 with that material next week.

• The first unit is going through a pre assembly of all its parts to verify bolt and part fit-up. Minor bolt length problems have appeared.

• MEI will glue the shims down to the pole tips as soon as the preassembly is complete.

• Cores for the Second batch of 6 UV Magnets remain in final machining.

Injector Dipoles (DU/DV)

• At Master Machine, GV cores & GU cores are being painted. Gap check on the first unit showed uniformity to .00014 inches.

Arc 180 Degree Dipoles (GY)

• The 2nd core is being shipped to Wang.

• At Wang NMR, the third GY Coil (of four) was started.

• GY Coil Potting fixture is in fabrication.

• GG Coils were potted.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) continued manufacturing the GQ-GX Cores.


3 inch quad (QX)

• We completed measuring all QX quads for both the IR and the UV machine.

Trim Quad (QT)

• Milhous Control of Virginia didn’t ship three magnets last week but claims to be shipping two today.

• Continuing the potential solution to the four required vertical correctors in the arcs. (GCs), Tom Hiatt's analyzed the case of add on coils in the corner that have the side coils as a partner rather than the top coil. It was successful.

Sextupole (SF)

• DULY Research continues final detailing of the sextupole.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The ARC Chambers are out for bid.

• The X Chamber remains in fabrication in the Shop.

• The drawings for the recirculation region in front of the Cryomodule will be signed off today.

• Design of the four chambers for the optical chicane region is now complete.

• Design of the regions around the wiggler continues.

WBS 11 (Optics):

We have been inspecting the OCMMS beamline parts, with some parts are at the machine shop being checked with the coordinate measuring machine. Tests of Ga:In eutectic with a narrower 0.005" gap were performed, and showed that the material still flowed. Since the In:Ag braze is working so well, and we can ill-afford to delay procuring the outcoupler holders, we are adopting this technique.

We've signed the drawings for the optical cavity vacuum vessels, and the procurement process is underway. We anticipate receipt by mid-August. The Lab machine shop says we are still on schedule (6/24) on delivery of a portion of the cavity internals. We received the 4" UHV gate valves that isolate the vacuum vessels from the rest of the beamline.

The ultrafast laser oscillator was installed this week. It is set up for 1 ps operation; next week the service engineer will return and set it up for ultrashort pulse operation. We have begun defining the research plan for damage testing UV optics with this laser.